# Markov Decision Processes                                  Under External Temporal Processes

### v2

@@ Abstract.
Most reinforcement learning algorithms treat the context under which they operate as a stationary, isolated, and undisturbed environment. However, in real world applications, environments constantly change due to a variety of external events. To address this problem, we study Markov Decision Processes (MDP) under the influence of an external temporal process. We formalize this notion and discuss conditions under which the problem becomes tractable with suitable solutions. We propose a policy iteration algorithm to solve this problem and theoretically analyze its performance. We derive results on the sample complexity of the algorithm and study its dependency on the extent of non-stationarity of the environment. We then conduct experiments to illustrate our results in a classic control environment.
@@ 1. Introduction
Reinforcement Learning (RL) provides a framework for sequential decision-making problems, wherein an agent has to make decisions at each time step so that a given objective is achieved. A huge amount of research has been done in this field in recent years, resulting in remarkable progress in various applications like robotics, finance, and even video games (Zhao et al., 2020; Charpentier et al., 2021; Mnih et al., 2015).
Central to RL is the mathematical framework of Markov Decision Processes (MDPs). In an MDP, an agent is assumed to be in an ‘environment’, and at each time step, the agent has to take a decision or an ‘action’. The state of the environment changes stochastically in a Markov manner. It depends only on the current state and the agent’s current action.
@@
While most RL research models the agent’s environment as an MDP, this paper argues that this is a very simplistic model of the real world, as one assumes that the state of a Markov Decision Process might constantly be changing, but the rules of the state transition are still fixed. However, in practical applications, external events may influence the agent’s environment and change its dynamics. For example, consider the problem of portfolio management in finance. An agent has access to a state consisting of various fundamental and technical indicators of a select number of financial instruments and has to make decisions, sequentially, on whether to hold on to an instrument, sell it, or short it, etc. But the way the prices of these instruments move may be affected by external forces, such as government decisions like rate hikes or sudden market movements caused by events occurring in a completely different sector of the economy not being considered by the agent. Such an external event might exert lasting influence over different time scales depending on the kind of event that occurred. A rate cut might affect stock prices for a few months, while a flash crash might cease its effect in a few hours.
@@
As another example, consider a navigation problem in which an agent has control of a robot and needs to travel from one point to another. While the agent can learn to do this task in ideal conditions, the optimal actions taken by the agent might depend on external conditions such as rain, sudden traffic changes, or human movement patterns. So, while an agent can learn to achieve goals in an ideal environment, it cannot be oblivious to external changes that influence the environment dynamics differently.
To tackle this problem, we consider a Markov Decision Process that is perturbed by an external discrete-time temporal process. This external process affects the dynamics of the MDP in a non-Markovian fashion, resulting in a non-stationary problem.
@@
Figure 1. An illustration of the lasting non-stationary effect of perturbations due to exogenous events. The plot shows the prices of two assets following a Geometric Brownian motion. A few exogenous Gaussian shocks to one asset forces it to diverge significantly from the other over the long term, with a markedly different pattern of evolution long after the effect of the exogenous event decays.
@@
    We formulate a non-stationary reinforcement learning problem consisting of an MDP perturbed by an external discrete-time temporal process that affects the dynamics of the MDP in a non-Markovian fashion. We describe sufficient conditions under which such a problem has a well-defined solution.
@@
    We derive conditions under which a solution can be determined approximately by an agent based on just the current state and a finite history of past events. This is possible when the perturbations caused by events older than $t$ time steps on the MDP transition dynamics and the event process itself are bounded in total variation by $M_{t}$ and $N_{t}$ respectively, and $\sum_{t}M_{t},\sum_{t}N_{t}$ are convergent series. We analyze the trade-off between the performance of the agent and the length of history considered.
@@
    We propose a policy iteration algorithm to solve this problem and theoretically analyze its behavior. We show that policy improvement occurs in regions of the state space where the Bellman error is not too low based on $M_{t}$ and $N_{t}$. Higher the approximation error due to non-stationarity, smaller the region where the policy is guaranteed to improve.
@@
2. Related Work
While most Reinforcement Learning research considers a stationary Markov Decision Process, there has been some work done on tackling various kinds of non-stationarities that might affect the MDP. Many works consider environments that are piece-wise stationary, i.e., the transition and reward functions remain unchanged except during a select few time steps called change points, during which they can change arbitrarily. Different works consider different assumptions on prior knowledge about the nature of the change points etc., available to the agent.
Padakandla et al. (2020) assume that the transition and reward functions belong to a fixed set of ‘contexts’ between change points. They propose a model-free algorithm that uses a change point detection algorithm that determines when a change occurs by analyzing the state-reward-next state tuples obtained from the environment. However, their algorithm assumes knowledge of the pattern of changes in the context.
Alegre et al. (2021) propose a model-based algorithm that can deal with arbitrarily different types of contexts and continuous state and action spaces. Other works consider drifting environments wherein the transition and reward functions can vary at any time step, while satisfying some constraints to make the problem tractable. Lecarpentier and Rachelson (2019) consider a finite horizon MDP in which the transition dynamics and reward function evolve such that they are Lipschitz continuous with respect to time. Cheung et al. (2020) consider such a setting in a finite MDP where the amount of total change in the environment, referred to as the variation budget, is known to the agent. They propose an optimistic Value Iteration-based algorithm and prove guarantees on the upper bound of its dynamic regret.
Wei and Luo (2021) propose a black-box reduction that turns any suitable algorithm with optimal performance in a near-stationary environment into an algorithm that achieves optimal dynamic regret in a non-stationary environment.
Hadoux et al. (2014) formalized a class of piecewise stationary MDPs called Hidden-Semi-Markov-Model MDPs where the transition from one hidden stationary MDP mode to another is semi-Markov in nature. To solve this problem, they adapt the Partially Observable Monte Carlo Planning (POMCP) algorithm used for solving Partially Observable Markov Decision Processes (POMDPs).
Some algorithms assume additional structure on the nature of non-stationarity. Feng et al. (2022) propose Factored Non-stationary MDPs (FN-MDPs), where the transition dynamics are defined in terms of a Dynamic Bayesian Network over the states, actions, rewards, and latent change factors that cause non-stationarity. Here, both the states and the latent change factors have transition dynamics obeying this causal graph. They propose using a Variational Auto-Encoder (VAE) framework to learn transition dynamics and present results in MuJoCo and robotic manipulation tasks.
Another factorization setting with a resemblance to ours has been proposed by Efroni et al. (2022) called the Exogenous MDP, wherein the environment state is divided into two parts, an endogenous state which is controllable by the agent’s actions, and an exogenous state, which evolves on its own and does not affect the agent’s reward. The precise way in which this division exists is unknown to the agent. While this may seem similar to our formulation of states and external events, in our case, the external events do influence the dynamics of the state and thereby, the rewards obtained by the agent.
A similar decomposition of the state space into endogenous and exogenous sub-spaces was studied by Dietterich et al. (2018), who assumed that the exogenous states evolve in a Markov fashion and the reward function can be additively decomposed into an exogenous reward and an endogenous reward. While their work studies how to decompose the given state into endogenous and exogenous components, in contrast, our work deals with essentially the opposite task, which is to study the effect of incorporating exogenous event information into reinforcement learning algorithms.
3. Preliminaries
@@
A Markov Decision Process (MDP) is defined by a tuple $(S,A,\left\{A(s):s\in S\right\},Q,c,\gamma)$, where $S$ is the state space, the set of all possible states of the environment, $A$ is the action space, the set of all possible actions that can be taken by the agent, $A(s)$ is the set of allowed actions at state $s$, $Q$ is the transition probability kernel, $c$ is the cost function, and $\gamma$ is the discount factor.
At each time step $t$, the agent is in state $s\in S$ and has to take action $a\in A(s)\subseteq A$. Depending on the action taken, a cost $c(s,a)$ is incurred by the agent, and the environment transitions to state $s^{\prime}\in S$ sampled from $Q\left(.|(s,a)\right)$, which is then communicated to the agent. This process repeats indefinitely.
In general, $S$ and $A$ are Borel spaces, $A(s)$ is a non-empty measurable subset of $A$ for each $s\in S$, $c$ is a measurable function, and $\gamma\in(0,1)$.
@@
$Q$ is a stochastic kernel on $S$ given $\mathbb{K}$, where $\mathbb{K}=\left\{(s,a):s\in S,a\in A(s)\right\}$ is the set of all feasible state-action pairs. That is, $Q(.|y)$ is a probability measure on $S$ for each $y\in\mathbb{K}$, and $Q(B|.)$ is a measurable function on $\mathbb{K}$ for each $B\in\mathcal{B}(S)$, the Borel $\sigma$-algebra on $S$.
Any rule $\pi=\{\pi_{t}\}$ by which the agent chooses the action, as a function of state and possibly time, is called a policy. A policy can be either deterministic with $\pi_{t}(s)=a_{t}\in A(s)$, or stochastic with $\pi_{t}(.|s\>)$ being a stochastic kernel on $A(s)$ given $s\in S$.
A policy that is independent of time is called a stationary policy. The goal of the agent is to determine a policy $\pi$ among all possible policies $\Pi$ such that the following performance criterion is minimized:
$V^{\pi}(s)=\mathsf{E}_{s_{t},a_{t}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],\qquad\pi\in\Pi,x\in S,$
the sum of discounted costs incurred by the agent when it starts from a state and follows policy $\pi$ at each time step, resulting in a trajectory of states $s_{0}=s,s_{1},s_{2},\dots,$ by taking actions $a_{0},a_{1},a_{2},\dots$. The function $V^{\pi}$ is called the value function of the policy $\pi$.
By the Monotone Convergence Theorem, the above can be written as the limit
$V^{\pi}(s)=\lim_{n\rightarrow\infty}V_{n}^{\pi}(s),$
where $V_{n}$ is the $n$-stage cost, the sum of discounted costs obtained from the first $n$ time steps.
@@
A policy $\pi^{*}$ that achieves the minimum value is called an optimal policy and the corresponding value function $V^{\pi^{*}}$ is denoted as $V^{*}$.
3.1. Bellman optimality equation
@@
A measurable function $v:S\rightarrow\mathbb{R}$ is said to be a solution to the Bellman optimality equation if it satisfies:
$v(s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
The right-hand side of this equation can be considered as an operator $\mathcal{T}$ acting on the function $v$, leading to the optimal Bellman equation
$v=\mathcal{T}v,$
where the optimal Bellman operator $\mathcal{T}$ is defined as
$[\mathcal{T}v](s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
A function that satisfies the optimal Bellman equation is a fixed point of the operator $\mathcal{T}$. Under suitable conditions, the optimal value function satisfies this equation.
Assumption 1.
  1. (1)
    The one-stage cost $c$ is lower semi-continuous, non-negative and inf-compact on $\mathbb{K}$.
  2. (2)
@@
Assumption 2.
There exists a policy $\pi$ such that $V^{\pi}(s)<\infty$ for all $s\in S$.
Under the above assumptions, we can make some statements about the existence and uniqueness of the optimal policy and value function, given by the following result.
Theorem 3.1 (Theorem 4.2.3 of Hernández-Lerma and Lasserre (2012)).
Suppose the above two assumptions hold, then:
  1. (1)
    The optimal value function $V^{*}$ is the point-wise minimal solution to the above Bellman equation. If $u(.)$ is any other solution to the above equation, then $u(.)\geq V^{*}(.)$.
  2. (2)
    There exists a function $\pi^{*}:S\rightarrow A$ such that $\pi^{*}(s)\in A(s)$ that attains the minimum in the Bellman equation, i.e.,
    $V^{*}(s)=c(x,\pi^{*}(s))+\gamma\int V^{*}(s^{\prime})Q(ds^{\prime}|s,\pi^{*}(s))\quad\text{for all }s\in S,$
    and the deterministic stationary policy $\pi^{*}$ is optimal. Conversely, any deterministic stationary policy that is optimal satisfies the above equation.
  3. (3)
    If an optimal policy exists, then there exists one that is deterministic and stationary.
4. MDP under an external temporal process
Consider an MDP $\mathcal{M}=(S,A,Q,r,\gamma)$ where
    $S$: The state space, a closed subset of $\mathbb{R}^{m}$
    $A$: The action space, a closed subset of $\mathbb{R}^{n}$
    $\mathbb{K}=S\times A$: The set of all feasible state-action pairs
    $r:S\times A\times S\rightarrow\mathbb{R}$: Reward function as a function of the state-action pair and the resultant next state
    $\gamma$: Discount factor
    $Q$: A stochastic kernel on $S$ given $\mathbb{K}$
We assume that all actions in $A$ are allowed at every state in $S$, so there is no need to specify the allowed action set separately. Also, instead of a cost function $c$, we consider a reward function $r$, where the problem objective is maximize the discounted sum of rewards instead.
Consider an external discrete-time temporal process that is independent of the above MDP but affects the transition probabilities of the MDP.
The probability of an event occurring in the external process at time $t$ is a function of all the previous (external) events till that time. Associated with each event is a probabilistic mark $x\in\mathcal{X}$, where $\mathcal{X}$ is a closed subset of $\mathbb{R}^{d}$.
Suppose at time $t$ that the state of the MDP is $s\in S$. Suppose the history of external events till that time is $H_{t}=\{(t_{1},x_{1}),(t_{2},x_{2}),\dots,(t_{j},x_{j})\}$, where the ordered pair $(t_{i},x_{i})$ denotes an event with mark $x_{i}$ that occurred at time $t_{i}$.
Then, this history perturbs the stochastic kernel $Q$ of the MDP to give rise to a new transition kernel $Q_{H_{t}}$ on $S$ given $\mathbb{K}$.
Similarly, this history also determines the next event that occurs in the temporal process. We denote the distribution of the event mark at time $t$ by $Q^{X}_{H_{t}}$, which is a probability distribution on $\mathcal{X}$. Since this describes the event distribution that is external to the MDP, it does not depend on the state of the MDP or the action taken by the agent.
4.1. Assumptions
4.1.1. On the external process
  1. (i)
    There exists an event with mark, say $X=0$ (zero), that is equivalent to a non-event.
  2. (ii)
    Old events have a reduced influence on the current probability transition kernel of the MDP. More specifically, there exists a convergent series $\sum_{T}M_{T}$ of real numbers, such that at any time $t$, for any event histories $H_{t}$ and $H^{\prime}_{t}$, if the MDP is in state $s$ and action $a$ is taken, and if $H_{t}$ and $H^{\prime}_{t}$ differ by only one event at time $t^{\prime}$, then
    $\mathsf{TV}\left(Q_{H_{t}}(.|s,a),Q_{H^{\prime}_{t}}(.|s,a)\right)<M_{T},\quad\text{for all }(s,a)\in\mathbb{K},$
  3. (iii)
    Similarly, old events also have a reduced influence on the probability distribution of the current event mark. Specifically, there exists a convergent series $\sum_{T}N_{T}$ of real numbers such that for event histories $H_{t}$ and $H^{\prime}_{t}$ at time $t$ that differ at only one event at time $t^{\prime}$, if $t-t^{\prime}\geq T$, then
    $\mathsf{TV}\left(Q^{X}_{H_{t}},Q^{X}_{H^{\prime}_{t}}\right)<N_{T}.$
4.1.2. Regularity
Define a new function $r_{P}:\mathbb{K}\rightarrow\mathbb{R}$ induced from $r$ by $P$ as
where $P$ is any stochastic kernel on $S$ given $\mathbb{K}$. We assume the following regularity conditions on $r$ and $Q$:
  1. (i)
  2. (ii)
  3. (iii)
    $r_{P}$ is sup-compact on $\mathbb{K}$, i.e., for every $s\in S,r\in\mathbb{R}$, and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process, the set $\{a\in A:r_{P}(s,a)\geq r\}(\subseteq A)$ is compact.
  4. (iv)
    $Q$ is strongly continuous.
  5. (v)
@@ 4.1.2. Regularity
Note: A transition kernel $P$ is said to be strongly continuous if
@@
is continuous and bounded on $\mathbb{K}$ for every measurable bounded function $v$ on $S$.

### v3

@@ Abstract.
Most reinforcement learning algorithms treat the context under which they operate as a stationary, isolated, and undisturbed environment. However, in real world applications, environments constantly change due to a variety of external events. To address this problem, we study Markov Decision Processes (MDP) under the influence of an external temporal process. First, we formalize this notion and derive conditions under which the problem becomes tractable with suitable solutions. We propose a policy iteration algorithm to solve this problem and theoretically analyze its performance. Our analysis addresses the non-stationarity present in the MDP as a result of non-Markovian events, necessitating the formulation of policies that are contingent upon both the current state and a history of prior events. Additionally, we derive insights regarding the sample complexity of the algorithm and incorporate factors that define the exogenous temporal process into the established bounds. Finally, we perform experiments to demonstrate our findings within a traditional control environment.
@@ 1. Introduction
Reinforcement Learning (RL) serves as a foundational framework for addressing sequential decision-making problems. In recent years, extensive research in this domain has led to significant advancements across various fields, including robotics, finance, and navigation. At the core of RL lies the mathematical framework of Markov Decision Processes (MDPs). In this framework, an agent interacts with an ‘environment’, where, at each time step, the agent is required to make a decision to take an ‘action.’ The state of the environment changes stochastically in accordance with the Markov property, relying solely on the present state and the action taken by the agent. While solving the various sequential decision-making problems, in most RL research, it is assumed that while the state of an MDP may be constantly changing, the rules governing state transitions remain unchanged. However, in practical applications, external events may influence the agent’s environment and change its dynamics.
@@
For example, in the context of portfolio management in finance, an agent has access to a state consisting of various fundamental and technical indicators of a select number of financial instruments and has to make decisions sequentially on whether to hold on to an instrument, sell it, short it, etc. However, the price movements of these instruments may be affected by external forces, such as government decisions like rate hikes or sudden market movements caused by events occurring in a completely different sector of the economy that are not being considered by the agent. Such an external event might exert lasting influence over different time scales depending on the kind of event that occurred. A rate cut might affect stock prices for a few months, while a flash crash might cease its effect in a few hours. An example illustrating the persistent non-stationarity effect of perturbations due to exogenous events is depicted in Figure 1.
@@
As another example, consider a navigation problem in which an agent has control of a robot and needs to travel from one point to another. While the agent can learn to do this task in ideal conditions, the optimal actions taken by the agent might depend on external conditions such as rain, sudden traffic changes, or human movement patterns. Hence, while an agent can learn to achieve goals in an ideal environment, it cannot be oblivious to external changes that influence the environment dynamics differently.
@@
Figure 1. The plot shows the prices of two assets following a Geometric Brownian motion. A few exogenous Gaussian shocks to one asset forces it to diverge significantly from the other over the long term, with a markedly different pattern of evolution long after the effect of the exogenous event decays.
In the literature, non-stationarity settings in RL have been studied, albeit considering only very special cases, mostly providing practical solutions. Some model non-stationarity as piece-wise stationarity (Padakandla et al., 2020; Alegre et al., 2021; Hadoux et al., 2014), in which the environment remains unchanged except during a select few time steps called change points, during which it can change arbitrarily. Few works consider drifting environments wherein the transition and reward functions can vary at any time step while satisfying some constraints to make the problem tractable (Lecarpentier and Rachelson, 2019; Cheung et al., 2020). Hallak et al. (2015) study Contextual MDPs, in which the environment dynamics change based on externally specified “contexts” that are episodic and possibly unknown chosen from among a finite set of known values. Tennenholtz et al. (2023) extend this setting to handle dynamic contexts that are known to the agent, change at each time step, and influence the state transition distribution. However, their contexts are not exogenous and come from a finite set of possible vectors.
In contrast to the above existing literature, we study this problem in a general setting, where a Markov Decision Process is perturbed by an external discrete-time temporal process. This external process affects the dynamics of the MDP in a non-Markovian fashion, resulting in a non-stationary problem. This process, being exogenous to the decision process, does not depend on the state of the environment under consideration or the actions of the agent. In this paper, we study conditions that the exogenous process has to satisfy, along with its effect on the MDP, such that almost optimal solutions exist that are tractable in some sense. We formulate the variants of standard reinforcement learning algorithms that consider above-mentioned non-stationarity problem and theoretically analyze the properties of these algorithms.
@@
    We present a non-stationary reinforcement learning formulation for an MDP under the influence of an external discrete-time temporal process, which alters the MDP dynamics in a non-Markovian manner. We outline the necessary conditions that ensure the existence of a well-defined solution for this problem.
@@
    We establish criteria under which an approximate solution can be determined using only the current state and finite history of past events. We show that this is possible when the perturbations caused by events older than $t$ time steps on the MDP transition dynamics and the event process itself are bounded in total variation by $M_{t}$ and $N_{t}$ respectively, and $\sum_{t}M_{t},\sum_{t}N_{t}$ are convergent series. We analyze the trade-off between the performance of the agent and the length of history considered.
@@
    We propose a policy iteration algorithm and theoretically analyze its behavior. We show that policy improvement occurs in regions of the state space where the Bellman error is “not too low” based on $M_{t}$ and $N_{t}$. Based on our analysis, we observe that the higher the approximation error due to non-stationarity, the smaller the region where the policy is guaranteed to improve.
@@
Organization
In Section 2, we provide mathematical details of MDPs in a continuous setting to provide a background on conditions for the existence and uniqueness of optimal policy and value function. A reader who is familiar with these concepts can skip this Section. In Section 3, we set the stage to solve the main problem, and we state the assumptions and regularity conditions. We derive guarantees for the existence of almost optimal solutions in Section 4, and a policy iteration algorithm in Section 5. We provide sample complexity analysis in Section 6 and experiments in Section 7. We provide a detailed literature survey and the position of our work in the existing literature in Section 8.
2. Preliminaries: Markov Decision Processes
@@
Consider a Markov Decision Process (MDP) defined by a tuple $\mathcal{M}=(S,A,Q,r,\gamma)$, where $S$ is the state space, the set of all possible states of the environment, $A$ is the action space, the set of all possible actions that can be taken by the agent, $Q$ is the transition probability kernel, $r$ is the reward function, and $\gamma$ is the discount factor. At each time step $t$, the agent is in state $s\in S$ and has to take action $a\in A$ obtaining reward $r(s,a)$, and the environment transitions to state $s^{\prime}\sim Q\left(\;.\;|s,a\right)$.
@@
In general, $S$ and $A$ are Borel spaces [^1] A complete separable metric space equipped with a Borel $\sigma$-algebra, $r$ is a measurable function, and $\gamma\in(0,1)$. $Q$ is a stochastic kernel on $S$ given $S\times A$. That is, $Q(.|y)$ is a probability measure on $S$ for each $y\in S\times A$, and $Q(B|.)$ is a measurable function on $S\times A$ for each $B\in\mathcal{B}(S)$, the Borel $\sigma$-algebra on $S$.
A policy is a rule used to choose an action based on the state. The goal of the agent is to determine a policy $\pi$ among all possible policies $\Pi$ that achieves the maximum possible value function $V^{\pi}$, which is defined as the expected discounted reward obtained by the agent following policy $\pi$:
$V^{\pi}(s)=\mathsf{E}_{s_{t},a_{t}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],\qquad\pi\in\Pi,s\in S,$
@@
where the expectation is over the states and actions in the agent’s trajectory due to the stochasticity of the environment and possibly the policy itself. A policy $\pi^{*}$ that achieves the optimal value is called an optimal policy, and its corresponding value function $V^{\pi^{*}}$ is denoted as $V^{*}$.
@@
A measurable function $v:S\rightarrow\mathbb{R}$ is said to be a solution to the Bellman optimality equation if it satisfies $v=\mathcal{T}v$, where $\mathcal{T}$ is the optimal Bellman operator defined as
$[\mathcal{T}v](s)=\max_{a\in A}\left[r(s,a)+\gamma\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
A function that satisfies the optimal Bellman equation is a fixed point of the operator $\mathcal{T}$. Under suitable regularity conditions on the rewards and transition kernel, the optimal value function $V^{*}$ satisfies this equation. Suppose the reward function $r$ is upper semi-continuous, bounded, and sup-compact, and the transition kernel $Q$ is strongly continuous, i.e., the function $v(s,a)=\mathsf{E}_{s^{\prime}\sim Q(.|s,a)}[v(s^{\prime})]$ is continuous and bounded for every measurable bounded function $v$ on $S$. Then, there exists a unique optimal policy $\pi^{*}$ whose value function $V^{*}$ is the solution of the Bellman optimality equation (Theorem 4.2.3 of Hernández-Lerma and Lasserre (1996)).
3. MDP under an External Temporal Process
Consider an continuous MDP $\mathcal{M}=(S,A,Q,r,\gamma)$ and an external discrete-time temporal process that is independent of the above MDP but affects the transition probabilities of the MDP. The probability of an event occurring in the external process at time $t$ is a function of all the past (external) events till that time. Associated with each event is a probabilistic mark $x\in\mathcal{X}$, where $\mathcal{X}$ is a closed subset of $\mathbb{R}^{d}$.
Suppose at time $t$ that the state of the MDP is $s\in S$. Denote the history of external events till time $t$ by $H_{t}=\{(t_{1},x_{1}),(t_{2},x_{2}),\dots,(t_{j},x_{j})\}$, where the ordered pair $(t_{i},x_{i})$ denotes an event with mark $x_{i}$ that occurred at time $t_{i}$. History $H_{t}$ perturbs the stochastic kernel $Q$ of the MDP to give rise to a new transition kernel $Q_{H_{t}}$ on $S$ given $S\times A$. At the same time, this history also determines the next event that occurs in the temporal process. We denote the distribution of the event mark at time $t$ by $Q^{X}_{H_{t}}$, which is a probability distribution on $\mathcal{X}$. Since this describes the event distribution that is external to the MDP, it does not depend on the state of the MDP or the action taken by the agent.
We assume the following.
  1. (A1)
    There exists an event with a mark, say $X=0$ (zero), that is equivalent to a non-event.
  2. (A2)
    Events that occurred in the distant past have a reduced influence on the current probability transition kernel of the MDP. More specifically, there exists a convergent series $\sum_{T}M_{T}$ of real numbers, such that at any time $t$, for any event histories $H_{t}$ and $H^{\prime}_{t}$, if the MDP is in state $s$ and action $a$ is taken, and if $H_{t}$ and $H^{\prime}_{t}$ differ by only one event at time $t^{\prime}$, then
    $\mathsf{TV}\left(Q_{H_{t}}(.|s,a),Q_{H^{\prime}_{t}}(.|s,a)\right)<M_{T},\quad\text{for all }s\in S,a\in A,$
  3. (A3)
    Similarly, events that occurred in the distant past also have a reduced influence on the probability distribution of the current event mark. Specifically, there exists a convergent series $\sum_{T}N_{T}$ of real numbers such that for event histories $H_{t}$ and $H^{\prime}_{t}$ at time $t$ that differ at only one event at time $t^{\prime}$, if $t-t^{\prime}\geq T$, then
    $\mathsf{TV}\left(Q^{X}_{H_{t}},Q^{X}_{H^{\prime}_{t}}\right)<N_{T}.$
Along with the above assumptions (A1-A3), the MDP $\mathcal{M}$ under the influence of the exogenous event process needs to satisfy some regularity conditions for guaranteeing the existence and uniqueness of optimal solutions. Define a new function $r_{P}:S\times A\rightarrow\mathbb{R}$ induced from $r$ by $P$ as
where $P$ is any stochastic kernel on $S$ given $S\times A$. We assume the following regularity conditions on $r$ and $Q$:
  1. (B1)
  2. (B2)
  3. (B3)
    $r_{P}$ is sup-compact on $S\times A$, i.e., for every $s\in S,r\in\mathbb{R}$, and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process, the set $\{a\in A:r_{P}(s,a)\geq r\}(\subseteq A)$ is compact.
  4. (B4)
@@ 3. MDP under an External Temporal Process
  5. (B5)
@@
Note that a transition kernel $P$ is said to be strongly continuous if
@@
is continuous and bounded on $S\times A$ for every measurable bounded function $v$ on $S$.